HIGH RESOLUTION INFRARED SPECTROSCOPY OF CO_{2}-Br_{2}

\underline{A. SAZONOV}, R. A. BEAUDET, Department of Chemistry, University of Southern California, Los Angeles, CA 90089-0482.

The absorption spectrum of the weakly bound complex CO_{2}-Br_{2} has been observed by probing the asymmetric stretch of the CO_{2} moiety near 2349 cm^{−1}. Complex was formed by supersonic expansion of the mixture of CO_{2} and bromine vapor with use He as a carrier gas. The CO_{2}-Br_{2} was found to have a linear structure with a Br atom close to the center of mass of the system. The isotope substitution of the other Br provided the splitting of the observed peaks into two. The measured rotational and distortional constants for ^{79}Br and ^{81}Br are B′ = 0.0148044(26), D′ = 6.43(17) × 10^{−8}, B′′ = 0.0146591(26), D′′ = 6.01(18) × 10^{−8}, and B′ = 0.0146765(27), D′ = 6.42(18) × 10^{−8}, B′′ = 0.0145322(28), D′′ = 6.03(19)×10^{−8}, respectively (in cm^{−1}). The experimental values will be compared with ab initio calculation results and the floppiness of the molecule will be discussed.